OpenStudy (anonymous):

#1: When you heat an air-filled balloon, what happens inside with regard to the movement of air molecules? A. The collisions of air molecules against the wall of the balloon are weaker. B. The collisions of air molecules against the wall of the balloon are stronger. C. The collisions of air molecules against the wall of the balloon do not change. **my answer: B is that right?? @aaronq :)

OpenStudy (anonymous):

#2: You have 0.5 L of air at 203 K in an expandable container at constant pressure. You heat the container to 273 K. What is the volume of air? A. 0.37 L B. 0.67 L C. 1.5 L D. 0.25 L **myanswer: B. 0.67 L is that right?

OpenStudy (anonymous):

#3: You place a balloon in a closed vacuum chamber at standard temperature and pressure (STP). You decrease the chamber temperature by half but keep the pressure constant. What happens to the balloon? A. It reduces in size by half. B, It stays the same volume. C. It doubles in volume. D. It expands to four times the original volume. *my answer: A is that irhgt?

OpenStudy (aaronq):

they both are correct

OpenStudy (aaronq):

3 is good too

OpenStudy (anonymous):

#4: A balloon contains 0.5 L of air at 273 K. Its volume increases to volume of 0.75 L at constant pressure. What is the temperature of air inside the balloon? A. 137 K B. 182 K C. 205 K D. 410 K **my answer: D. 410 K is that right?? and yayy :)

OpenStudy (aaronq):

for #4, if you increase the volume the temp will go down

OpenStudy (anonymous):

#5: This question appears on a quiz. "A 1.0 L balloon has a temperature of 273 K. When the temperature changes to 30°C, what is the volume?" A student answers 0.11 L. Which of the following explanations is true? (C = K 273; K = C + 273.) A. The student did not solve for the correct parameter, and the answer should be 1.11 L. B. The student did not solve for the correct parameter, and the answer should be 0.90 L. C. The student did not put the temperatures in the same units, and the answer should be 0.90 L. D. The student did not put the temperatures in the same units, and the answer should be 1.11 L. **My answer: D not too sure tho... is that right? :/

OpenStudy (anonymous):

so #4 is wrong then>?? idk what it is :(

OpenStudy (aaronq):

for #4, you use P1V1=P2V2 plug the numbers in and you can't ever get a wrong answer.. unless you put them in the wrong place

OpenStudy (anonymous):

i don't even know haha :/ like how do u plug them in? cuz i definitely didn't do it right :/

OpenStudy (aaronq):

for #5, the "student" divided 30/273.. so i would say wrong parameters.. idk how someone would do that though

OpenStudy (anonymous):

like what goes where?

OpenStudy (anonymous):

wait so the answer to #5 is this? A. The student did not solve for the correct parameter, and the answer should be 1.11 L.

OpenStudy (aaronq):

P=pressure V= volume on one side is the set for the system at one point, then across the equal sign for the change

OpenStudy (aaronq):

for #5, 1.11L is not the correct answer...

OpenStudy (anonymous):

so #5 is B. The student did not solve for the correct parameter, and the answer should be 0.90 L. ???? and I'm confused about #4 now haha :/ could u pls show me how u do it? :/

OpenStudy (anonymous):

heres how i think it is based off what u said.. :/ .5/273=.75/x and solve for x? is that right?

OpenStudy (anonymous):

and i got x=409.5 which is about 410K.. which is D.. but you said that #4 isn't D right?? :/ so i don't know what to do now :( @aaronq what do i do for it then?

OpenStudy (aaronq):

no they're not being divided

OpenStudy (aaronq):

PV=nRT it's basically based off the ideal gas law, PV/nRT of one system equals PV/nRT of the other \[\frac{ P _{1}V _{1} }{ n _{1}RT _{1} }=\frac{ P _{2}V _{2} }{ n _{2}RT _{2} }\] when all the variables like T or n are constant, it simplifies to P1V1=P2V2

OpenStudy (anonymous):

okay so how do i apply that with the values that they give me?

OpenStudy (aaronq):

so for number 4, (0.5 L)(273 K)= (0.75 L)(T) (0.5 L)(273 K)/(0.75 L) = T

OpenStudy (anonymous):

oh so i get 182.. which is answer choice B right?

OpenStudy (anonymous):

and was #5's answer B. The student did not solve for the correct parameter, and the answer should be 0.90 L. ??

OpenStudy (aaronq):

yes

OpenStudy (anonymous):

yes to both 4 and 5? haha

OpenStudy (aaronq):

4 yes, 5 idk the student did both, not change his temperature units, and divided the wrong way...

OpenStudy (anonymous):

okay... so which would u go with for 5?? cuz idk haha

OpenStudy (aaronq):

i would go with C, because the student still got an answer in litres... somehow lol

OpenStudy (anonymous):

okay haha :) thanks!!!:)

OpenStudy (aaronq):

no prob! :)